OpenStudy (anonymous):

what are the critical points of 30(x-2)^2(x-4)^3

OpenStudy (anonymous):

Differentiate that! 30 * (2(x-2)*(x-4)^3 + 3(x-4)^2(x-2)) Set equal to zero... 30 * (2(x-2)*(x-4)^3 + 3(x-4)^2(x-2)) = 0 2(x-2)*(x-4)^3 + 3(x-4)^2(x-2)) = 0 x = 4 and x = 2

OpenStudy (anonymous):

I'd love a "become a fan" button clicked :D

OpenStudy (anonymous):

sorry i meant the critical points of f'(x) = 30(x-2)^2(x-4)^3

OpenStudy (anonymous):

So of that function...? Then you gotta' differentiate it anyways and you'd get the same thing.